OpenStudy (anonymous):

Joanne's Dress Shop received an invoice dated July 25 for $1,400, with terms of 2/10, 1/15, n/60. On August 8, Joanne's Dress shop sent a partial payment of $750. The actual amount that should be credited is? please help me i need the answer asap

OpenStudy (mertsj):

We don't know what you mean when you say terms of 2/10, 1/15 and so on.

OpenStudy (anonymous):

@Mertsj the 2/10 and so on wouldbe considered fractions

OpenStudy (tkhunny):

2% Discount if paid in 10 days. 1% Discount if paid in 15 days, (but greater that 10 days) Net due in 60 days. 1400*0.98 = 1372 will close the acount on days 1-10 1400*0.99 = 1386 will close the acount on days 11-15 August 8 is Day 15 and falls under the 1% discount rule. 750/0.99 = 757.58 from the balance, and 1400.00 - 757.58 = 642.42 is due by day 60.

OpenStudy (tkhunny):

Let's do another one, only you show your talents!

OpenStudy (anonymous):

thamk you @tkhunny